# Memory-Augmented Quantum Reservoir Computing

### v1

@@ Abstract
Reservoir computing (RC) is an effective method for predicting chaotic systems by using a high-dimensional dynamic reservoir with fixed internal weights, while keeping the learning phase linear, which simplifies training and reduces computational complexity compared to fully trained recurrent neural networks (RNNs). Quantum reservoir computing (QRC) uses the exponential growth of Hilbert spaces in quantum systems, allowing for greater information processing, memory capacity, and computational power. However, the original QRC proposal requires coherent injection of inputs multiple times, complicating practical implementation. We present a hybrid quantum-classical approach that implements memory through classical post-processing of quantum measurements. This approach avoids the need for multiple coherent input injections and is evaluated on benchmark tasks, including the chaotic Mackey-Glass time series prediction. We tested our model on two physical platforms: a fully connected Ising model and a Rydberg atom array. The optimized model demonstrates promising predictive capabilities, achieving a higher number of steps compared to previously reported approaches.
@@ I Introduction
Reservoir Computing (RC) is a computational framework that has gained significant attention for efficiently processing time-dependent data. Originally inspired by recurrent neural networks (RNNs), RC exploits the dynamics of a fixed, high-dimensional system, known as the reservoir, to transform input signals into a rich, dynamic representation [1, 2]. This transformation allows for the linear extraction of complex features, making RC particularly suitable for tasks such as time series prediction, pattern recognition, and control systems. The key advantage of RC lies in its simplicity and efficiency. Unlike traditional RNNs, where the entire network is trained, RC requires training only on the output layer, while the reservoir remains untrained[1, 2, 3, 4, 5, 6, 7]. RC systems can be implemented in quantum domains, exploiting quantum mechanics properties to enhance computational power and memory capacity further. Quantum computing offers new paradigms in information processing by utilizing principles like superposition, entanglement, and quantum interference, allowing for exponential increases in computational power and memory space. These features enable quantum systems to tackle problems that are computationally infeasible for classical systems [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. The first proposal of Quantum Reservoir Computing (QRC) by Fujii and Nakajima [20] utilizes the space of quantum density matrices to store and process information, by injecting the input data into the state of one qubit of a fully connected qubit chain.
@@
Building on that foundational work, researchers have investigated a variety of platforms, each offering specific advantages for Quantum Reservoir Computing (QRC) [21, 22, 23, 24]. Among these, spin chains have been extensively studied across a wide range of tasks, from time series prediction to Extreme Learning Machines [25, 26, 27, 28, 29, 30, 31, 32]. Additionally, fermionic and bosonic systems have been proposed as alternative platforms [33, 34, 35]. Other explored platforms include quantum oscillators [36, 37], photonic systems [38], and Rydberg atoms, which offer strong dipole-dipole interactions as a key feature [39]. Furthermore, gate-based platforms have recently been explored for their universality and high controllability in implementing QRC algorithms [40, 41, 42, 43, 44, 45, 46, 47].
@@
Despite their theoretical promise, protocols based on Fujii’s approach require multiple copies of the quantum system, rendering them experimentally impractical. To address this challenge, hybrid approaches have been proposed [43, 34, 47, 48], drawing inspiration from classical reservoir computing [49, 50, 51, 52]. In these hybrid models, quantum evolution is combined with classical post-processing, providing a more feasible solution for practical implementation. However, these hybrid systems often suffer from a significant drawback: the loss of memory concerning previous steps. Feedback mechanisms, wherein the result of the previous step is reintroduced into the system at the subsequent step, have been proposed to address this issue [43, 32, 48].
@@
In our work, we develop a quantum-classical hybrid approach, incorporating a straightforward classical post-processing component that enhances memory capacity to fully utilize classical memory resources. We evaluate our algorithm on several standard tasks, including linear memory, NARMA, and, ultimately, a time series prediction task: the Mackey-Glass series prediction. We evaluate our approach using two distinct physical systems: a fully connected Ising model in a transverse field and a platform based on Rydberg atoms.
@@
Figure 1: Schematic representation of the proposed hybrid quantum-classical reservoir computing (QRC) method. The input data $\{s_{k}\}$ is injected into the quantum reservoir by modulating a parameter of the Hamiltonian $H_{k}$, which governs the dynamics of a quantum system (e.g., a transverse-field Ising model or Rydberg atom array). The system evolves for a time $\Delta t$, during which the quantum state $|\psi_{k}(\Delta t)\rangle$ is influenced by the input. After the time evolution, the output is obtained by measuring a set of observables $\vec{m}_{k}$, which carry information about the input data. These measurements are further processed classically to construct a reservoir state $\vec{r}_{k}$ that retains the memory of previous inputs. A linear fit is then applied to the reservoir state to produce the predicted output, thereby allowing the system to solve complex tasks such as time series prediction with enhanced memory capacity.
@@ II Method
We introduce a model of quantum reservoir computing where the input is injected into the system by varying a Hamiltonian parameter. Our model uses a Hamiltonian similar to that proposed by Fukii and Nakajima [20], specifically a transverse-field Ising model with an additional input-dependent magnetic field aligned with the coupling terms:
@@
where $J_{ij}$ represents the coupling constants, $h$ is the static magnetic field, and $h_{k}$ is the additional magnetic field containing the information about the $k$-th value of the input sequence $\{s_{k}\}$. For each input $s_{k}$ a target function is identified $y_{k}(\{s_{k^{\prime}}\}_{k^{\prime}\leq k})$ that generically depends on the previous values of the input sequence.
For each input data, a pure state |$\ket{\psi_{0}}$ is evolved for a $\Delta t$ according to the Hamiltonian $\hat{H}_{k}$, |$\ket{\psi_{0}^{k}(\Delta t)}=e^{-i\hat{H}_{k}\Delta t}\ket{\psi_{0}}$. After the time evolution, the expectation values of a set of observables, $\vec{m_{k}}$, are measured.
@@
$\hat{S}^{n_{S}}$ and $\hat{B}^{l_{r}}$, act respectively on the reservoir state at the previous step and the measurement vector, and $\gamma$ is a real-valued parameter between 00 and $1$ that weighs the previous input. $\hat{S}^{n_{S}}$ acts on a generic vector by shifting the vector elements by $n_{S}$ steps forward, if positive, or backward, if negative, (${\left[\hat{S}^{n_{S}}\right]}_{ij}=\delta_{\left(i+n_{S}\right)\bmod l_{r},j}$). It coincides with a cyclic permutation of the reservoir vector and can be represented by an orthogonal matrix. Its role is to enhance the distinguishability of the inputs at different times by avoiding the same observable being summed up at different times, thus enhancing classical memory. The operator $\hat{B}^{l_{r}}$ acts on the measurement vector by lengthening it, interleaving zeros between the elements such that its final length is $l_{r}$, which coincides with the length of $\vec{r_{k}}$. As an example, given a vector $\{v_{1},v_{2},v_{3},v_{4}\}$, the action of the operator $\hat{S^{1}}$ coincides with $\hat{S}^{1}\{v_{1},v_{2},v_{3},v_{4}\}=\{v_{2},v_{3},v_{4},v_{1}\}$, while applying $\hat{B}^{8}$ introduces four zeros, resulting in $\hat{B}^{8}\{v_{1},v_{2},v_{3},v_{4}\}=\{v_{1},0,v_{2},0,v_{3},0,v_{4},0\}$
@@
To simulate the Reservoir Computing algorithm, we perform exact diagonalization of $N$=$5$ spins, with $J_{ij}$ randomly generated from a uniform distribution in $[-J,J]$, $h$=$0.2J$, $h_{k}$=$C_{s}s_{k}$, $C_{s}=0.2$, $\gamma=0.95$ for the memory tasks and $\gamma=0.6$ for the prediction task. All energies and times are presented respectively in units of $J$ and $1/J$. A table with names for the sets of observables is provided below in Tab. 1.
@@ III.1 STM task
To begin our analysis of the performance of our Reservoir Computing (RC) algorithm, we examine the simplest task: assessing the system’s short-term memory (STM) capabilities. This task assesses the reservoir’s capability to store and retrieve past input sequences. In this task, the input sequence $\{s_{k}\}$ is randomly generated from a uniform distribution over the interval $[-1,1]$. The goal is to obtain the input value from $\tau$ time steps earlier: $y_{k}=s_{k-\tau}$.
@@
As a comparison, we also display the case without the cyclic permutation operator ($n_{s}=0$), red line in Fig. 2, where $\Delta t$ has been set to its best value ($\Delta t=0.01$). It is evident that the cyclic permutation operator is crucial for retaining significant memory: the red line is almost zero for $\tau>2$.
@@
Figure 2: Top) Memory Capacity (C) for the STM task as a function of $\tau$ for several $\Delta t$ values of the Hamiltonian evolution. The set of observables $\vec{m}^{A}$ is used. Bottom) Memory Capacity as a function of $\tau$ for several reservoir lengths $l_{r}$ determined by the $\hat{B}^{l_{r}}$ operator when only one observable is used ($\vec{m}^{X1}$). In both panels, the cyclic permutation operator is implemented with $n_{s}=1$, except for the red line, for which $n_{s}=0$.
@@ III.2 NARMA task
To further check the performance of our RC algorithm, we test it for the Nonlinear Auto-Regressive Moving Average (NARMA) task. The NARMA task is a benchmark used to evaluate the non-linear memory and computational capabilities of RC systems. It involves generating a sequence where the next value depends non-linearly on the previous values and a random input sequence. The NARMA task with order $n$ is defined as:
@@
The performance on the NARMA task is measured by how accurately the reservoir can reproduce the target output sequence, again we use the square of the Pearson coefficient ($C$).
In Fig. 3 (top), we demonstrate that, in contrast to the linear memory task, the performance on the NARMA task does not reach its peak for the smallest $\Delta t$, but for $\Delta t\approx 0.05$. This is because the NARMA task requires some degree of non-linearity, provided by the time evolution, to achieve optimal performance. Also in this case, the non-linear memory can be significantly enhanced by increasing the reservoir dimension using the lengthening operator $B^{n_{B}}$, as illustrated in Fig. 3 (bottom).
Figure 3: Top) C for the NARMA task as a function of the NARMA order $n$ for several $\Delta t$ values of the Hamiltonian evolution. The set of observables $\vec{m}^{A}$ is used. Bottom) C as a function of $n$ for several reservoir lengths $l_{r}$, determined by the $\hat{B}^{l_{r}}$ operator when only one observable is used ($\vec{m}^{X1}$). In both panels, the cyclic permutation operator is implemented with $n_{s}=1$.
In Fig. 4, we examine the non-linear memory capacity for the NARMA task using different sets of observables, while keeping the reservoir dimension fixed at ${l_{r}45}$. It is observed that increasing the number of observables does not lead to performance improvements. The best results, as in the STM case, are achieved with a single observable ($\vec{m}^{X1}$), which is experimentally more friendly, as it simplifies the implementation and reduces the complexity of the measurement process. This suggests that for the NARMA task, a single observable is sufficient to capture the necessary information for optimal performance. In this case, measuring more observables leads to a small decrease in performance. The reason is that the reservoir dimension is fixed, and increasing the number of observables leaves less room for classical memory.
@@ III.3 Prediction task
where $\epsilon=0.3$ is a predefined error tolerance and $\sigma$ is the standard deviation of the time series.
@@
From Fig. 5, it is evident that the cyclic permutation function is crucial, as the prediction capability is negligible without it, and there are slight performance differences between various permutation functions, i.e., $n_{s}$=±$\pm 1,\pm 2$. The optimal evolution time for the spin network is $\Delta t\approx 10$ in units of $J$, significantly longer than the time required for memory tasks. This indicates that more complex tasks necessitate greater non-linear processing and signal distribution across the network nodes. The extended evolution time allows the system to better integrate and process the temporal dependencies inherent in the MG equation, thereby enhancing prediction accuracy.
Figure 6: Top) VPT for the MG series prediction as a function of the reservoir dimension ł$\l_{r}$ for different sets of observables. For each point, the mean value and standard deviation of the mean are shown, based on $30$ different realizations of the Hamiltonian. Bottom) Example of the predicted time series for the best reservoir dimension ($l_{r}=315$).
Fixing $\Delta t$ at its optimal value, Fig. 6 (top) shows the Valid Prediction Time (VPT) as a function of the reservoir dimension, adjustable via the operator $\hat{B}^{l_{r}}$, for different sets of observables. Firstly, a general improvement in VPT is observed with increasing the reservoir dimension for all the sets of observables, up to a saturation point around $270$. Using different sets of observables produces slightly varying results. Good performance is achieved with commutating observables ($\vec{m}^{XX}$) or single-particle observables ($\vec{m}^{N}$), while non-commutating and double-spin observables do not offer benefits ($\vec{m}^{1-N}$). Conversely, using the biggest considered set of observables ($\vec{m}^{A}$) provides a slight improvement in VPT. In Fig. 6 (bottom), we present the true and predicted MG series for a specific case, with the set of observables $\vec{m}^{A}$. It is seen that the system can predict up to approximately 700 steps, which is significantly higher than recent literature results [26, 43, 48].

### v2

# Memory-Augmented Hybrid Quantum Reservoir Computing

@@ Abstract
Reservoir computing (RC) is an effective method for predicting chaotic systems by using a high-dimensional dynamic reservoir with fixed internal weights, while keeping the learning phase linear, which simplifies training and reduces computational complexity compared to fully trained recurrent neural networks (RNNs). Quantum reservoir computing (QRC) uses the exponential growth of Hilbert spaces in quantum systems, allowing for greater information processing, memory capacity, and computational power. We present a hybrid quantum-classical approach that implements memory through classical post-processing of quantum measurements, thus, avoiding the need for multiple coherent input injections (as in the original QRC proposal). We tested our model on two physical platforms: a fully connected Ising model and a Rydberg atom array, and evaluated it on various benchmark tasks, including the chaotic Mackey-Glass time series prediction, where it demonstrates significantly enhanced predictive capabilities and achieves a substantially longer prediction time, outperforming previously reported approaches.
@@ I Introduction
Reservoir Computing (RC) is a computational framework that has gained significant attention for efficiently processing time-dependent data. Originally inspired by recurrent neural networks (RNNs), RC exploits the dynamics of a fixed, high-dimensional system, known as the reservoir, to transform input signals into a rich, dynamic representation [1, 2]. This transformation allows for the linear extraction of complex features, making RC particularly suitable for tasks such as time series prediction, pattern recognition, and control systems. The key advantage of RC lies in its simplicity and efficiency. Unlike traditional RNNs, where the entire network is trained, RC requires training only on the output layer, while the reservoir remains untrained[1, 2, 3, 4, 5, 6, 7]. RC systems can be implemented in quantum domains, exploiting quantum mechanics properties to further enhance computational power and memory capacity. Quantum computing offers new paradigms in information processing by utilizing principles like superposition, entanglement, and quantum interference, allowing for exponential increases in computational power and memory space. These features enable quantum systems to tackle problems that are computationally infeasible for classical systems [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. The first proposal of Quantum Reservoir Computing (QRC) by Fujii and Nakajima [20] utilizes the space of quantum density matrices to store and process information, by injecting the input data into the state of one qubit of a fully connected qubit network.
@@
Building on that foundational work, researchers have investigated a variety of platforms, each offering specific advantages for Quantum Reservoir Computing (QRC) [21, 22, 23, 24]. Among these, spin chains have been extensively studied across a wide range of tasks, from time series prediction to Extreme Learning Machines [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41]. Additionally, fermionic and bosonic systems have been proposed as alternative physical implementation of reservoir [42, 43, 44]. Other explored set-ups include quantum oscillators [45, 46], photonic systemsfor both Quantum RC [47], and Quantum Extreme Learning Machines [48], and Rydberg atoms, which offer strong dipole-dipole interactions as a key feature [49, 50], with platforms such as Bloqade by QuEra enabling the simulation of many-body quantum systems. Furthermore, gate-based platforms have recently been explored for implementing QRC algorithms, due to their universality and high controllability [51, 52, 53, 54, 55, 56, 57, 58].
@@
Despite their theoretical promise, protocols based on Fujii’s approach require multiple copies of the quantum system, rendering them experimentally impractical. To address this challenge, hybrid approaches have been proposed [54, 43, 58, 59], drawing inspiration from classical reservoir computing [60, 61, 62, 63]. In these hybrid models, quantum evolution is combined with classical post-processing, providing a more feasible solution for practical implementation. However, these hybrid systems often suffer from a significant drawback: the loss of memory concerning previous steps. Feedback mechanisms, wherein the result of the previous step is reintroduced into the system at the subsequent step, have been proposed to address this issue [54, 33, 59, 64].
@@
In this work, we introduce a novel quantum-classical hybrid approach that integrates a simple classical post-processing component to boost memory capacity by exploiting classical memory resources, bridging between QRC and Extreme Learning Machine. We assess the performance of our algorithm across several benchmark tasks, including linear memory, NARMA, and, finally, time series prediction using the Mackey-Glass series. Our approach is tested on two different physical systems: a fully connected Ising model in a transverse field and a system utilizing Rydberg atoms. For both of the cases, we find a substantial enhancement of predictive capabilities.
@@
Figure 1: Schematic representation of the proposed hybrid quantum-classical reservoir computing (QRC) method. At each step $k$, the input data $\{s_{k}\}$ is encoded into the quantum reservoir by modulating a parameter of the Hamiltonian $H_{k}$, which generates the dynamics of a quantum system (e.g., a transverse-field Ising model or Rydberg atom array). The system evolves for a time $\Delta t$, giving rise to the quantum state $|\psi_{k}(\Delta t)\rangle$, depending on the input. After the time evolution, the output is obtained by measuring a set of observables $\vec{m}_{k}$, which carry information about the input data. These measurements are further processed classically to construct a reservoir state $\vec{r}_{k}$ that retains the memory of previous inputs. A linear fit is then applied to the reservoir state to produce the predicted output, thereby allowing the system to solve complex tasks such as time series prediction with enhanced memory capacity.
@@ II Method
We introduce a model of quantum reservoir computing where the input is encoded into the system by varying a Hamiltonian parameter. Our model uses a Hamiltonian similar to that proposed by Fujii and Nakajima [20]; namely, a transverse-field Ising model with an additional input-dependent magnetic field aligned with the coupling terms:
@@
where $J_{ij}$ represents the coupling constants, $h$ is the static magnetic field, and $h_{k}$ is the additional magnetic field containing the information about the $k$-th value of the input sequence $\{s_{k}\}$. For each input $s_{k}$ a target function $y_{k}(\{s_{k^{\prime}}\}_{k^{\prime}\leq k})$ is identified, which generically depends on the previous values of the input sequence.
For each input data, a pure state |$\ket{\psi_{0}}$ is evolved for a time $\Delta t$ according to the Hamiltonian $\hat{H}_{k}$, |$\ket{\psi_{0}^{k}(\Delta t)}=e^{-i\hat{H}_{k}\Delta t}\ket{\psi_{0}}$. After this evolution, the expectation values of a set of observables, $\vec{m_{k}}$, are measured.
@@
$\hat{S}^{n_{S}}$ and $\hat{B}^{l_{r}}$ act on the reservoir state at the previous step and the measurement vector, respectively, with $\gamma$ being a real-valued parameter between 00 and $1$ that weighs the previous input. $\hat{S}^{n_{S}}$ acts on a generic vector by shifting the vector elements by $n_{S}$ steps forward, if positive, or backward, if negative, (${\left[\hat{S}^{n_{S}}\right]}_{ij}=\delta_{\left(i+n_{S}\right)\bmod l_{r},j}$). It coincides with a cyclic permutation of the reservoir vector and can be represented by an orthogonal matrix. Its role is to enhance the distinguishability of the inputs at different times by avoiding the same observable being summed up at different times, thus boosting classical memory. The operator $\hat{B}^{l_{r}}$ acts on the measurement vector by lengthening it, interleaving zeros between the elements such that its final length is $l_{r}$, which coincides with the length of $\vec{r_{k}}$. As an example, given a vector $\{v_{1},v_{2},v_{3},v_{4}\}$, the action of the operator $\hat{S^{1}}$ coincides with $\hat{S}^{1}\{v_{1},v_{2},v_{3},v_{4}\}=\{v_{2},v_{3},v_{4},v_{1}\}$, while applying $\hat{B}^{8}$ introduces four zeros, resulting in $\hat{B}^{8}\{v_{1},v_{2},v_{3},v_{4}\}=\{v_{1},0,v_{2},0,v_{3},0,v_{4},0\}$.
@@
To simulate the Reservoir Computing algorithm, we perform exact diagonalization of $N$=$5$ spins, with $J_{ij}$ randomly generated from a uniform distribution in $[-J,J]$, $h$=$0.2J$, $h_{k}$=$C_{s}s_{k}$, $C_{s}=0.2$, $\gamma=0.95$ for the memory tasks and $\gamma=0.6$ for the prediction task. All energies and times are presented in units of $J$ and $1/J$, respectively. A table with names for the sets of observables is provided below in Tab. 1.
@@ III.1 STM task
To begin the analysis of the performance of our Reservoir Computing (RC) algorithm, we examine the simplest task: assessing the system’s short-term memory (STM) capabilities, which measures the reservoir’s ability to store and retrieve past input sequences. In this task, the input sequence $\{s_{k}\}$ is randomly generated from a uniform distribution over the interval $[-1,1]$. The goal is to obtain the input value from $\tau$ time steps earlier: $y_{k}=s_{k-\tau}$.
@@
As a comparison, we also display the case without the cyclic permutation operator ($n_{s}=0$), red line in Fig. 2, where $\Delta t$ has been set to its best value ($J\Delta t=0.01$). It is evident that the cyclic permutation operator is crucial for retaining significant memory: the red line is almost zero for $\tau>2$.
@@
Figure 2: Top) Memory Capacity (C) for the STM task as a function of $\tau$ for several $\Delta t$ values of the Hamiltonian evolution. The set of observables $\vec{m}^{A}$ is used. The cyclic permutation operator is employed in all cases (with $n_{s}=1$), except for the red line, for which $n_{s}=0$. Bottom) Memory Capacity as a function of $\tau$ for several reservoir lengths $l_{r}$ determined by the $\hat{B}^{l_{r}}$ operator when only one observable is used ($\vec{m}^{X1}$).
@@ III.2 NARMA task
To further check the performance of our RC algorithm, we test it for the Nonlinear Auto-Regressive Moving Average (NARMA) task, which is a benchmark usually employed to evaluate the non-linear memory and computational capabilities of RC systems. It involves generating a sequence where the next value depends non-linearly on both the previous ones and a random input sequence. The NARMA task with order $n$ is defined as:
@@
The performance on the NARMA task is measured by the accuracy with which the reservoir reproduces the target output sequence; to quantify it, we use again the square of the Pearson coefficient ($C$).
In Fig. 3 (top), we demonstrate that, in contrast to the linear memory task, the performance on the NARMA task does not reach its peak for the smallest $\Delta t$, but for $J\Delta t\approx 0.05$. This is because the NARMA task requires some degree of non-linearity, provided by the time evolution, to achieve optimal performance. In this case too, the non-linear memory can be significantly enhanced by increasing the reservoir dimension using the lengthening operator $B^{n_{B}}$, as illustrated in Fig. 3 (bottom).
Figure 3: Top) Pearson coefficient $C$ for the NARMA task, as a function of the NARMA order $n$ for several $\Delta t$ values of the Hamiltonian evolution. The set of observables $\vec{m}^{A}$ is used. Bottom) C as a function of $n$ for several reservoir lengths $l_{r}$, determined by the $\hat{B}^{l_{r}}$ operator when only one observable is used ($\vec{m}^{X1}$). In both panels, the cyclic permutation operator is implemented with $n_{s}=1$.
In Fig. 4, we examine the non-linear memory capacity for the NARMA task using different sets of observables, while keeping the reservoir dimension fixed at ${l_{r}45}$. It is observed that increasing the number of observables does not lead to performance improvements. The best results, as in the STM case, are achieved with a single observable ($\vec{m}^{X1}$), which is experimentally more friendly, as it simplifies the implementation and reduces the complexity of the measurement process. This suggests that for the NARMA task, a single observable is sufficient to capture the necessary information for optimal performance. In this case, measuring more observables leads to a small decrease in performance. This is due to the fact that we fixed the reservoir dimension, so that increasing the number of observables leaves less room for classical memory.
@@ III.3 Prediction task
where $\epsilon=0.3$ is a predefined error tolerance and $\sigma$ is the standard deviation of the time series. We express VPT in integer units of the discretized time.
@@
From Fig. 5, it is evident that the cyclic permutation function is crucial, as the prediction capability is negligible without it, and there are slight performance differences between various permutation functions, i.e., $n_{s}$=±$\pm 1,\pm 2$. The optimal evolution time for the spin network is $J\Delta t\approx 10$, significantly longer than the time required for memory tasks. This indicates that more complex tasks necessitate greater non-linear processing and signal distribution across the network nodes. The extended evolution time allows the system to better integrate and process the temporal dependencies inherent in the MG equation, thereby enhancing prediction accuracy.
Figure 6: Top) VPT for the MG series prediction as a function of the reservoir dimension ł$\l_{r}$ for different sets of observables. For each point, the mean value and standard deviation of the mean are shown, based on $30$ different realizations of the Hamiltonian. Bottom) Example of the predicted time series for the best reservoir dimension ($l_{r}=315$) as a function of the integer discretized time.
Setting $\Delta t$ at its optimal value, Fig. 6 (top) shows the Valid Prediction Time (VPT) as a function of the reservoir dimension, adjustable via the operator $\hat{B}^{l_{r}}$, for different sets of observables. Firstly, a general improvement in VPT is observed with increasing the reservoir dimension for all the sets of observables, up to a saturation point around $270$. Using different sets of observables produces slightly varying results. Good performance is achieved with commutating observables ($\vec{m}^{XX}$) or single-particle observables ($\vec{m}^{N}$), while non-commutating and double-spin observables do not offer benefits ($\vec{m}^{1-N}$). Conversely, using the biggest considered set of observables ($\vec{m}^{A}$) provides a slight improvement in VPT. In Fig. 6 (bottom), we present the true and predicted MG series for a specific case, with the set of observables $\vec{m}^{A}$. It is seen that the system can predict up to approximately 700 steps, which is significantly higher than recent literature results [26, 54, 59].